# ALMA-IMF XV: NH kinematic analysis on the intermediate protocluster G353.41 Key Words.: stars: formation – ISM: clouds – ISM: kinematics and dynamics – ISM: molecules1 Introduction

### v1

@@ 4 NH isolated component velocity decomposition
In Fig. 3 we see that clear multiple isolated velocity components are present in our dataset. To characterize the complex dense-gas kinematics traced by NH we follow the method in Álvarez-Gutiérrez et al. (2021), and we use the spectroscopic toolkit PySpecKit (Ginsburg & Mirocha, 2011; Ginsburg et al., 2022) to model and decompose the isolated component emission. PySpecKit adjusts a fixed number of components set by the user, based on visual inspection of the data we impose three velocity components to every spectra and then remove false positives (see below). Given the kinematic complexity of the data and cursory inspection of the spectra, a simpler analysis with only two components contradicts the data. In essence, three components is the simplest possible choice, given the data. While this might fail for a small number of spectra that could require velocity components, the residuals indicate that this could occur in a severe minority of cases, and hence more components is not warranted given the SNR and resolution of this particular data set. To improve the convergence of PySpecKit, we create a set of ranges for the parameters that define each of the three Gaussian velocity components, namely the peak intensity, central velocity, and velocity dispersion. After testing different parameter ranges, we set the intensity range between 1.76 K (4 times the mean RMS) and 30 K, the velocity centroid from 30 km s to 20 km s, and the velocity dispersion from 0.22 km s to 1 km s.
@@
From the results using the ranges defined above, we notice that some modeled components do not fit any emission. These fits are the result of imposing to the fitter a fixed number of components, given these spectra can be better represented by one or two velocity components. In these fits there is no uncertainty estimation for both the peak intensity and velocity dispersion. Based on these two criteria we remove those velocity fits from the modeled cube. With this cleaning approach we are left with spectra characterized by one (, two (, and three ( Gaussian velocity components. We present the Gaussian fits of the high SNR spectra from Fig. 3 in Fig. B.1.
@@
In Fig. 5 we show the spatial distribution of the multiple Gaussian velocity components. In gray we indicate the main filamentary structure in the region (see § 3). The first and second velocity components, in blue and green respectively, present most of the high intensity emission and they also spatially dominate over the third, most red-shifted component. Both the first and second components trace mostly the filaments F1 and F3 from Fig. 4, where most of the 1.3 mm cores are located. The position of these cores coincide with high integrated intensity regions in these isolated velocity components. The most redshifted component is compact and less intense compared to the first and second velocity components. This velocity distribution is located mostly along the filament F2 and the central hub (see Fig. 4). In Fig. 6 we present the number of Gaussian velocity components for each spectra, where we highlight that:
Most of the NH data presents emission characterized by two velocity components.
@@
Most of the cores (black ellipses; Louvet et al. submitted) are located in regions with spectra presenting two to three Gaussian velocity components, indicating kinematic complexity even at 4 kau (NH spatial resolution).
@@
In Fig. 7 we show the histogram of the fitted velocity centroid of each Gaussian velocity component. The peaks of these distributions are located at 27, 24.7, and 23.3 km s respectively, well-separated in velocity. From hereafter we refer to these distribution as blue, green, and red respectively. Most of the velocity components appear to be associated with the blue and green distributions. For consistency with the different tracers used in further analysis, we shift the isolated component velocities by +8 km s, to the reference frame of the main line components of NH (Cazzoli et al., 1985).
4.1 DCN & NH derived core velocities
In this section our goal is to increase the sample of core velocities from the already published DCN catalog, aiming to explore all the potential in these types of dataset. Given the relatively high of DCN (32) (cm) compared to NH (10) (), DCN (32) is known to coincide well with continuum peaks associated to cores (Liu et al., 2015; Cunningham et al., 2016; Minh et al., 2018), while NH is characterized by tracing the dense gas at the innermost parts of star forming regions (Fernández-López et al., 2014; Hacar et al., 2018; González Lobos & Stutz, 2019).
@@ 4.1 DCN & NH derived core velocities
We use the ALMA-IMF DCN 12 m data from Cunningham et al. (2023), which presents a velocity resolution of km s. For each DCN velocity core described by a single component (Cunningham et al., 2023) we compare the emission of the DCN and modeled NH isolated spectra. We find an average velocity offset between the DCN peak and the closest NH isolated component peak of km s, less than two DCN channel widths. We use this approach for the remaining 30 cores, in order to determine their DCN velocities.
Here we estimate the RMS of the DCN data in emission-free channels in the range of 42 km s to 25 km s and we obtain the SNR map by dividing the peak intensity by the RMS. For the procedure below we only use DCN spectra with SNR 3. Next, we extract the average DCN and modeled NH isolated component spectra of these 30 cores. We identify the NH isolated velocity component closest to the DCN peak within three DCN channel widths. We find that 11 out of these 30 cores present DCN with SNR close to one NH velocity component. Here, we define the velocity of these cores as the velocity where the DCN emission peaks. On average, these cores have a velocity offset between these two tracers less than 0.8 km s ( DCN channels), similar to the results obtained for the 15 cores with DCN velocities from Cunningham et al. (2023), and they present an average velocity offset of 0.38 km s. Throughout this paper we refer to these cores as “DCN & NH cores” given they are derived from the comparison of these two tracers. In Table C.1 we include these DCN & NH core velocities, complementing the DCN catalogue from Cunningham et al. (2023). In Fig. C.1 we present two examples of DCN and NH spectra in cores where we see clear agreement between these tracers.
@@ 5.1 Traditional PV diagram
We start by analyzing the “traditional” PV diagram shown in Fig. 8. We create this diagram by taking the total intensity along the Galactic longitude, where indicates the distance in parsec relative to the center of G353, assuming a distance to the protocluster of of 2 kpc (Motte et al., 2022). We see general agreement between the DCN core velocities and the NH velocity distribution. This suggests that most of the cores are still kinematically coupled to the dense gas in which they formed. As presented in § 4.1, the DCN and NH velocities match within 0.8 km s ( DCN channels).
@@
Regarding the dense gas velocity distribution, in Fig. 8 we see a velocity spread of km s in the sub-region between 0.3 pc to 0.1 pc. Most of the intensity on this diagram is located at the upper part of this sub-region, at . This spread is also present in the PV diagram along and shown in the top right and bottom left panel of Fig. 9. We explore the possible origin of this structure in § 6.
@@ 5.2 Intensity-weighted PV diagrams
In the top left panel of Fig. 9 we show the spatial distribution of the fitted Gaussian velocity components (see § 4). The blue, green, and red color maps indicate the integrated intensity of the first, second, and third velocity components of the NH spectra respectively. Note that the spatial overlap between any of these components is presented in Fig. 6.
As seen in Fig. 8, the traditional PV diagram provides information on the dynamics on the large, protocluster-scale, environment. Meanwhile, the intensity-weighted position-velocity diagram (Fig. 9), where the color of each point indicates its integrated intensity, highlights the small core-scale kinematics. Similarly as in González Lobos & Stutz (2019) and Álvarez-Gutiérrez et al. (2021), from the isolated component line decomposition (§ 4), we derive the integrated intensity and velocity centroid for each Gaussian velocity component. Using these parameters we create intensity-weighted PV diagrams along the and coordinates. We present these NH PV diagrams in the bottom left and top right panels of Fig. 9. The key features on the position-position (PP) and on the top right PV diagram, are:
The agreement between the DCN core velocities and the overall NH PV structures suggests that cores are still kinematically coupled to the dense gas in which they formed.
We see at least nine clear and prominent V-shaped velocity gradients (see Fig. D.1), across all velocity components. The orientation of these V-shape, pointing to the left/right (top right panel) or up/down (bottom left panel), follow no clear preference.
@@
The most prominent V-shape is located at pc, 20.5 km s), between two 1.3 mm cores with DCN detections (see § 6).
@@ 5.3 Velocity gradients
In this section we focus on the most prominent blue V-shape (Fig. 9, top right panel). In Fig. 10 we show this velocity distribution in detail. In order to characterize the VGs composing this V-shape, we apply a linear fit to both the upper and lower VG. Given the visual linearity of the VGs composing the V-shape, we apply a linear fit to these distribution in order to characterize them. For these fits we consider data only above an integrated intensity threshold of 8 K km s and 3 K km s, for the upper and lower gradient respectively. We remove data not related to the velocity gradient, clustered in the ranges of , which lie just outside the filament hosting this V-shape on the POS. Additionally, we weight each point based on their integrated intensity to make our fits more robust. The slopes of the linear fits represent the VGs in km s pc. These linear fits follow the VGs distribution and these are somewhat asymmetric, the upper gradient is slightly shallower than the bottom gradient. Given the unknown inclination angle () of these structures relative to the POS, the observed VG is just a fraction of the original VG. These are related as VG = VG. These VGs present values between to km s pc (see Fig. 10). Additionally, we estimate the center of this V-shape as the velocity-weighted mean position of the points composing this structure. With this approach the position of the points closest to the V-shape apex present more weight, obtaining the center of this V-shape at () = (353.4135, 0.3657). This position is located between “core 2” and “core 3”, both of them having DCN velocity fits (Cunningham et al., 2023). These core present masses of 20.7 and 6.4 M respectively. We inspect the core catalog derived from the map at native resolution (Louvet et al. submitted) and the location of this V-shape do not coincide with any core.
@@
In the left panel of Fig. 11 we show the integrated intensity of the multiple modeled NH isolated components. With colored boxes we show the areas where we create the different PV diagrams presented on the right panel. These boxes are centered at the main blue V-shape, matching the area of this V-shaped structure (see Fig. 12). We show that the overall structure in PV space is conserved at different angles, excluding the possibility of this velocity feature being the result of projection effects.
@@
For this V-shaped structure, we note that the NH velocity range is V 2 km s, while the whole NH isolated component dataset present a V 8 km s (Fig. 9). From these VGs we derive timescales as = 1/VG, similar to the procedure for a rotating filament presented in Álvarez-Gutiérrez et al. (2021). We obtain timescales between kyr. These timescales are short compared to the 0.21 Myr free fall time (t) of the protocluster (Motte et al., 2022), and a few times larger than the t of nearby cores ( 20 kyr, within 0.1 pc of this V-shape). To determine the cores t, we use the 1.3 mm core masses from Louvet et al. (submitted). We characterized eight more NH V-shaped structure. These are presented in Appendix D. Note that within a beam size from the apex of V-shape “B” (see Fig. D.2), the continuum core “7” (M) is located.
@@
In Henshaw et al. (2014), they propose two scenarios that might produce these V-shaped velocity gradients (see their Fig. 12). One scenario suggests that gas in a filament is flowing towards a denser region (infall), while the other scenario suggests that a protostellar outflow moves the dense gas located in its vicinity. To analyze the different dynamical processes present in this region, we use the ALMA-IMF 12 m data of the shock, outflow tracer SiO (54), from Cunningham et al. (2023). From those data we create its intensity-weighted PV diagram, presented in Fig. E.1 (see Appendix E for more details). We note that there is almost no SiO emission nor outflow sources at the location of the NH V-shape (see Fig. 12). The V within 10 from this main blue V-shape is 40 km s, while for the whole SiO dataset is 80 km s, showing a clear difference in the SiO V and the core velocities. Furthermore, the SiO V is times larger than the one of NH. This difference in traced velocities implies that these two molecules trace vastly different physical phenomena. We suggest that the small velocity range probed by NH indicates that the velocity gradients can be considered as infall signatures (see § 6). We discuss the possible morphology of the filaments hosting V-shapes in § 8.
@@ 6 G353 as a collapsing region
We use the SiO () and CO () data from Towner et al. (2024) and Cunningham et al. (2023) to identify possible outflows near the V-shape presented in Fig. 10. For CO we measure the RMS in the emission-free velocity range of 145 to 286 km s. We use only CO data with SNR3 for our analysis. The cleaning of the SiO data is described in Appendix E. In Fig. 12, with a fuchsia “” we indicate the center of the V-shape from § 5.3, located between two 1.3 mm cores 2 & 3 (black ellipses), which present DCN velocity detections. From these diagrams we see there is neither SiO nor CO outflow detection at the location of the main blue V-shape (, ) = (0 pc, 0 pc). In the right panel we show the position of the data composing this V-shape, where the velocity peaks towards the center of this velocity feature.
We derive the mass-weighted mean position of the two cores (2 & 3) closest to the V-shape to determine their barycenter (yellow “” in Fig. 12). We find that the difference between the center of the V-shape and the barycenter of these cores is ″(au), well below the beam size of the NH data. A similar offset is also present in the intensity and velocity profiles along filaments from ATOMS data (Zhou et al., 2022, see their Fig. 6). This small spatial offset might suggest that the gas flowing in the V-shape is produced by the gravitational pull towards the barycenter of cores 2 & 3, where the NH radial velocities peak. This interpretation is similar to the one proposed in Zhou et al. (2023) in their kinematic analysis of the G333 complex. They describe the V-shaped velocity gradients as the result of gas funneling from the molecular cloud to clumps which is then funneled into cores (see their Fig. 9) consistent with gravitational acceleration.
In Fig. 13 we show the mean spectra of different tracers at the central position of the main blue V-shape. These spectra are measured over a circular region with diameter equal to the major axis of the NH beam (2.28 ″; ). This circular area results in a coverage of 1.14 times the NH beam, and 5.6 times the beams of the HCO, DCN, and HCO data. We see NH and HCO present double component spectra with asymmetric peaks. Between these peaks we detect DCN and HCO emission. The asymmetric spectra present in NH and HCO is consistent with the “blue asymmetry” spectral feature, usually interpreted as infall signature, suggesting that this region is under gravitational collapse (e.g. Anglada et al., 1987; Mardones et al., 1997; Lee et al., 1999, 2001; Smith et al., 2012). Based on the idea that the V-shapes are the result of flowing gas along filaments towards denser regions, the blue-asymmetry detected at the center of the main blue V-shape suggests that gravitational collapse is taking place at the apex of the V-shaped structure.
@@
For this comparison we model a sphere with a total mass of 150 M, a radius of 0.5 pc, and a power law density profile described by:
@@
where represents the horizontal coordinate in the POS, while represents the (non-observed) depth of the sphere.
In Fig. 14, we show the coverage of the PV distribution from our model with a solid white line. We find good agreement between the PV distributions of our approach and the data. The PV distribution of our infall model is consistent with previous work that provide the expected PV distributions for spherical protostellar envelopes under infall (Tobin et al., 2012). At large scales we interpret the agreement between the PV diagrams of our model and the data as protocluster scale collapse due to gravitational contraction. It is worth noting that the mass of our model 5.5 times lower than the mass derived from the (H) map (Díaz-González et al., 2023). We speculate that a model considering complex processes such as turbulence, radiative transfer, and magnetic fields might solve this mass discrepancy while still matching the observed PV distribution.

### v2

@@ 4 N2H+ isolated component velocity decomposition
In Fig. 3 we see that clear multiple isolated velocity components are present in our dataset. To characterize the complex dense-gas kinematics traced by N2H+ we follow the method in Álvarez-Gutiérrez et al. (2021), and we use the spectroscopic toolkit PySpecKit (Ginsburg & Mirocha, 2011; Ginsburg et al., 2022) to model and decompose the isolated component emission. PySpecKit adjusts a fixed number of components set by the user, based on visual inspection of the data we impose three velocity components to every spectrum and then remove false positives (see below). Given the kinematic complexity of the data and cursory inspection of the spectra, a simpler analysis with only two components contradicts the data. In essence, three components is the simplest possible choice, given the data. While this might fail for a small number of spectra that could require velocity components, the residuals indicate that this could occur in a severe minority of cases, and hence more components is not warranted given the SNR and resolution of this particular data set. To improve the convergence of PySpecKit, we create a set of ranges for the parameters that define each of the three Gaussian velocity components, namely the peak intensity, central velocity, and velocity dispersion. After testing different parameter ranges, we set the intensity range between 1.76 K (4 times the mean RMS) and 30 K, the velocity centroid from 30 km s-1 to 20 km s-1, and the velocity dispersion from 0.22 km s-1 to 1 km s-1.
@@
From the results using the ranges defined above, we notice that some modeled components do not fit any emission. These fits are the result of imposing to the fitter a fixed number of components, given these spectra can be better represented by one or two velocity components. In these fits there is no uncertainty estimation for both the peak intensity and velocity dispersion. Based on these two criteria we remove those velocity fits from the modeled cube. With this cleaning approach we are left with spectra characterized by one (, two (, and three ( Gaussian velocity components. We present the Gaussian fits of the high SNR spectra from Fig. 3 in Appendix B.
@@
In Fig. 5 we show the spatial distribution of the multiple Gaussian velocity components. In gray we indicate the main filamentary structure in the region (see § 3.1). The first and second velocity components, in blue and green respectively, present most of the high intensity emission and they also spatially dominate over the third, most red-shifted component. Both the first and second components trace mostly the filaments F1 and F3 from Fig. 4, where most of the 1.3 mm cores are located. The position of these cores coincide with high integrated intensity regions in these isolated velocity components. The most redshifted component is compact and less intense compared to the first and second velocity components. This velocity distribution is located mostly along the filament F2 and the central hub (see Fig. 4). In Fig. 6 we present the number of Gaussian velocity components for each spectrum, where we highlight that:
Most of the N2H+ data presents emission characterized by two velocity components.
@@
Most of the cores (black ellipses; Louvet et al. submitted) are located in regions with spectra presenting two to three Gaussian velocity components, indicating kinematic complexity even at 4 kau (N2H+ spatial resolution).
@@
In Fig. 7 we show the histogram of the fitted velocity centroid of each Gaussian velocity component. The peaks of these distributions are located at 27, 24.7, and 23.3 km s-1 respectively, well-separated in velocity. From hereafter we refer to these distribution as blue, green, and red respectively. Most of the velocity components appear to be associated with the blue and green distributions. For consistency with the different tracers used in further analysis, we shift the isolated component velocities by +8 km s-1, to the reference frame of the main line components of N2H+ (Cazzoli et al., 1985).
4.1 DCN & N2H+ derived core velocities
In this section our goal is to increase the sample of core velocities from the already published DCN catalog, aiming to explore all the potential in these types of dataset. Given the relatively high of DCN (32) (cm-3) compared to N2H+ (10) (), DCN (32) is known to coincide well with continuum peaks associated to cores (Liu et al., 2015; Cunningham et al., 2016; Minh et al., 2018), while N2H+ is characterized by tracing the dense gas at the innermost parts of star forming regions (Fernández-López et al., 2014; Hacar et al., 2018; González Lobos & Stutz, 2019).
@@ 4.1 DCN & N2H+ derived core velocities
We use the ALMA-IMF DCN 12 m data from Cunningham et al. (2023), which presents a velocity resolution of km s-1. For each DCN core velocity described by a single component (Cunningham et al., 2023) we compare the emission of the DCN and modeled N2H+ isolated spectra. We find an average velocity offset between the DCN peak and the closest N2H+ isolated component peak of km s-1, less than two DCN channel widths. We use this approach for the remaining 30 cores, in order to determine their DCN velocities.
Here we estimate the RMS of the DCN data in emission-free channels in the range of 42 km s-1 to 25 km s-1 and we obtain the SNR map by dividing the peak intensity by the RMS. For the procedure below we only use DCN spectra with SNR 3. Next, we extract the average DCN and modeled N2H+ isolated component spectrum of these 30 cores. We identify the N2H+ isolated velocity component closest to the DCN peak within three DCN channel widths. We find that 11 out of these 30 cores present DCN with SNR close to one N2H+ velocity component. Here, we define the velocity of these cores as the velocity where the DCN emission peaks. On average, these cores have a velocity offset between these two tracers less than 0.8 km s-1 ( DCN channels), similar to the results obtained for the 15 cores with DCN velocities from Cunningham et al. (2023), and they present an average velocity offset of 0.38 km s-1. Throughout this paper we refer to these cores as “DCN & N2H+ cores” given they are derived from the comparison of these two tracers. In Appendix C we present two examples of the comparison of the DCN and N2H+ spectra in cores where we see clear agreement between these tracers (Fig. C.1). In Table C.1 we show these obtained core velocities from our comparison between DCN & N2H+ spectra, complementing the DCN catalogue from Cunningham et al. (2023).
@@ 5.1 Traditional PV diagram
We start by analyzing the “traditional” PV diagram shown in Fig. 8. We create this diagram by taking the total intensity along the Galactic longitude, where indicates the distance in parsec relative to the center of G353, assuming a distance to the protocluster of of 2 kpc (Motte et al., 2022). We see general agreement between the DCN core velocities and the N2H+ velocity distribution. This suggests that most of the cores are still kinematically coupled to the dense gas in which they formed. As presented in § 4.1, the DCN and N2H+ velocities match within 0.8 km s-1 ( DCN channels).
@@
Regarding the dense gas velocity distribution, in Fig. 8 we see a velocity spread of km s-1 in the sub-region between 0.3 pc to 0.1 pc. Most of the intensity on this diagram is located at the upper part of this sub-region, at . This spread is also present in the PV diagram along and shown in the top right and bottom left panel of Fig. 9. We explore the possible origin of this structure in § 6.
@@ 5.2 Intensity-weighted PV diagrams
In the top left panel of Fig. 9 we show the spatial distribution of the fitted Gaussian velocity components (see § 4). The blue, green, and red color maps indicate the integrated intensity of the first, second, and third velocity components of the N2H+ spectra respectively. Note that the spatial overlap between any of these components is presented in Fig. 6.
As seen in Fig. 8, the traditional PV diagram provides information on the dynamics on the large, protocluster-scale, environment. Meanwhile, the intensity-weighted position-velocity diagram (Fig. 9), where the color of each point indicates its integrated intensity, highlights the small core-scale kinematics. Similarly as in González Lobos & Stutz (2019) and Álvarez-Gutiérrez et al. (2021), from the isolated component line decomposition (§ 4), we derive the integrated intensity and velocity centroid for each Gaussian velocity component. Using these parameters we create intensity-weighted PV diagrams along the and coordinates. We present these N2H+ PV diagrams in the bottom left and top right panels of Fig. 9. The key features on the position-position (PP) and on the top right PV diagram, are:
The agreement between the DCN core velocities and the overall N2H+ PV structures suggests that cores are still kinematically coupled to the dense gas in which they formed.
We see at least nine clear and prominent V-shaped velocity gradients (see Appendix D), across all velocity components. The orientation of these V-shape, pointing to the left/right (top right panel) or up/down (bottom left panel), follow no clear preference.
@@
The most prominent V-shape is located at pc, 20.5 km s-1), between two 1.3 mm cores with DCN detections (see § 6).
@@ 5.3 Velocity gradients
In this section we focus on the most prominent blue V-shape (Fig. 9, top right panel). In Fig. 10 we show this velocity distribution in detail. In order to characterize the VGs composing this V-shape, we apply a linear fit to both the upper and lower VG. Given the visual linearity of the VGs composing the V-shape, we apply a linear fit to these distribution in order to characterize them. For these fits we consider data only above an integrated intensity threshold of 8 K km s-1 and 3 K km s-1, for the upper and lower gradient respectively. We remove data not related to the velocity gradient, clustered in the ranges of , which lie just outside the filament hosting this V-shape on the POS. Additionally, we weight each point based on their integrated intensity to make our fits more robust. The slopes of the linear fits represent the VGs in km s-1 pc-1. These linear fits follow the VGs distribution and these are somewhat asymmetric, the upper gradient is slightly shallower than the bottom gradient. Given the unknown inclination angle () of these structures relative to the POS, the observed VG is just a fraction of the original VG. These are related as VG = VG. These VGs present values between to km s-1 pc-1 (see Fig. 10). Additionally, we estimate the center of this V-shape as the velocity-weighted mean position of the points composing this structure. With this approach the position of the points closest to the V-shape apex present more weight, obtaining the center of this V-shape at () = (353.4135∘, 0.3657∘). This position is located between “core 2” and “core 3”, both of them having DCN velocity fits (Cunningham et al., 2023). These core present masses of 20.7 and 6.4 M⊙ respectively. We inspect the core catalog derived from the map at native resolution (Louvet et al. submitted) and the location of this V-shape do not coincide with any core.
@@
In the left panel of Fig. 11 we show the integrated intensity of the multiple modeled N2H+ isolated components. With colored boxes we show the areas where we create the different PV diagrams presented on the right panel. These boxes are centered at the main blue V-shape, matching the area of this V-shaped structure (see Fig. 12). We show that the overall structure in PV space is conserved at different angles, excluding the possibility of this velocity feature being the result of projection effects.
@@
For this V-shape we use its composing VGs to derive timescales as = 1/VG, similar to the procedure for a rotating filament presented in Álvarez-Gutiérrez et al. (2021), and we suggest these can be interpreted as inflow timescales. The values for this V-shape range between kyr. These timescales are short compared to the 0.21 Myr free fall time (tff) of the protocluster (Motte et al., 2022), and a few times larger than the tff of nearby cores ( 20 kyr, within 0.1 pc of this V-shape). To determine the cores tff, we use the 1.3 mm core masses from Louvet et al. (submitted).
We characterized eight more N2H+ V-shaped structures. In Fig. 13 we show the position of these V-shapes in the POS. Previous studies have detected velocity gradients along filaments, towards a converging point (Peretto et al., 2014; Pan et al., 2024; Rawat et al., 2024). In the case of G353 we see that instead of detecting a single V-shape at the intersection of its filaments, the hub appears to be fragmented into multiple, small-scale V-shaped VG. Only V-shapes “A”, “D”, and “E” are outside of the hub, with V-shape “D” located on top of filament “F3”. This indicates that the V-shaped structures are not exclusive to the central parts of HFS, but are also present in comparatively isolated regions. Note that within a beam size from the apex of V-shape “B” (see Fig. D.2), the continuum core “7” (M⊙) is located.
@@
In Henshaw et al. (2014), they propose two scenarios that might produce these V-shaped velocity gradients (see their Fig. 12). One scenario suggests that gas in a filament is flowing towards a denser region (infall), while the other scenario suggests that a protostellar outflow moves the dense gas located in its vicinity. To analyze the different dynamical processes present in this region, we use the ALMA-IMF 12 m data of the shock, outflow tracer SiO (54), from Cunningham et al. (2023). From those data we create its intensity-weighted PV diagram, presented in Fig. E.1 (see Appendix E for more details). We note that there is almost no SiO emission nor outflow sources at the location of the N2H+ V-shape (see Fig. 12). The V within 10′′ from this main blue V-shape is 40 km s-1, while for the whole SiO dataset is 80 km s-1, showing a clear difference in the SiO V and the core velocities. Furthermore, the SiO V is times larger than the one of N2H+. This difference in traced velocities implies that these two molecules trace vastly different physical phenomena. We suggest that the small velocity range probed by N2H+ indicates that the velocity gradients can be considered as infall signatures (see § 6). We discuss the possible morphology of the filaments hosting V-shapes in § 8.
@@ 6 G353 as a collapsing region
We use the SiO () and CO () data from Towner et al. (2024) and Cunningham et al. (2023) to identify possible outflows near the V-shape presented in Fig. 10. For CO we measure the RMS in the emission-free velocity range of 145 to 286 km s-1. We use only CO data with SNR3 for our analysis. The cleaning of the SiO data is described in Appendix E. In Fig. 12, with a fuchsia “” we indicate the center of the V-shape from § 5.3, located between two 1.3 mm cores 2 & 3 (black ellipses), which present DCN velocity detections. From these diagrams we see there is neither SiO nor CO outflow detection at the location of the main blue V-shape (, ) = (0 pc, 0 pc). In the right panel we show the position of the data composing this V-shape, where the velocity peaks towards the center of this velocity feature.
We derive the mass-weighted mean position of the two cores (2 & 3) closest to the V-shape to determine their barycenter (yellow “” in Fig. 12). We find that the difference between the center of the V-shape and the barycenter of these cores is ″(au), well below the beam size of the N2H+ data. A similar offset is also present in the intensity and velocity profiles along filaments from ATOMS data (Zhou et al., 2022, see their Fig. 6). This small spatial offset might suggest that the gas flowing in the V-shape is produced by the gravitational pull towards the barycenter of cores 2 & 3, where the N2H+ radial velocities peak. This interpretation is similar to the one proposed in Zhou et al. (2023) in their kinematic analysis of the G333 complex. They describe the V-shaped velocity gradients as the result of gas funneling from the molecular cloud to clumps which is then funneled into cores (see their Fig. 9) consistent with gravitational acceleration.
In Fig. 14 we show the mean spectrum of different tracers at the central position of the main blue V-shape. These spectra are measured over a circular region with diameter equal to the major axis of the N2H+ beam (2.28 ″; ). This circular area results in a coverage of 1.14 times the N2H+ beam, and 5.6 times the beams of the H2CO, DCN, and H213CO data. We see N2H+ and H2CO present double component spectrum with asymmetric peaks. Between these peaks we detect DCN and H213CO emission. The asymmetric spectrum present in N2H+ and H2CO is consistent with the “blue asymmetry” spectral feature, usually interpreted as infall signature, suggesting that this region is under gravitational collapse (e.g. Anglada et al., 1987; Mardones et al., 1997; Lee et al., 1999, 2001; Smith et al., 2012). Based on the idea that the V-shapes are the result of flowing gas along filaments towards denser regions, the blue-asymmetry detected at the center of the main blue V-shape suggests that gravitational collapse is taking place at the apex of the V-shaped structure.
@@
For this comparison we model a sphere with a total mass of 150 M⊙, a radius of 0.5 pc, and a power law density profile described by:
@@
where represents the horizontal coordinate in the POS, while represents the (non-observed) depth of the sphere. The spatial coordinates for this model range from pc to pc.
In Fig. 15, we show the coverage of the PV distribution from our model, described in Eqs. 2 4, with a solid white line. We find good agreement between the PV distributions of our approach and the data. The PV distribution of our infall model is consistent with previous work that provide the expected PV distributions for spherical protostellar envelopes under infall (Tobin et al., 2012). At large scales we interpret the agreement between the PV diagrams of our model and the data as protocluster scale collapse due to gravitational contraction. It is worth noting that the inferred mass from our model is 5.5 times lower than the one derived from the (H2) map (Díaz-González et al., 2023). We speculate that a model considering complex processes such as turbulence, radiative transfer, and magnetic fields might solve this mass discrepancy while still matching the observed PV distribution.